# A Comment on the Higher-Spin Gauge Models

### v1

@@ Abstract
By applying the BRST deformation scheme[5] in the BRST-antifield formalism, higher-spin gauge models of massless totally-symmetric tensors are constructed in [6]222 See [7, 8, 9, 10, 11] for other attempts by using BRST cohomological metod. . Subsequently, these are generalized to include massless totally-symmetric tensor-spinors in [12]. For example, the higher-spin gauge model of two bosons $\{\phi,\ \phi_{1}\}$ is given as
@@
where $g$ is the deformation parameter. $S^{0}$ represents the free action while $S^{n}$ ($n\geq 1$) does the ($n+2$)-point vertex composed of $n+2$ (anti-)fields, given by333 See [12] for the notations used throughout this paper and the construction method of our models.
$S^{0}\ =$ $\int d^{D}x\ \text{tr}\Bigg{[}\frac{1}{2}\phi_{1}^{T}G(\phi_{1})+\phi^{*}_{1}{}^{T}\partial c+\frac{1}{2}\phi^{T}G(\phi)+\phi^{*}{}^{T}\partial c\Bigg{]},$ (2)
$S^{n}\ =$ $\int d^{D}x\ \text{tr}\Bigg{[}\frac{1}{2}\phi^{T}_{1}\Phi^{n}[G(\phi_{1})]+(\tilde{\partial}c\phi^{*}_{1})^{T}\Phi^{n-1}[G(\phi_{1})]+\frac{1}{2}(\tilde{\partial}c\phi^{*}_{1})^{T}\Phi^{n-2}[\tilde{G}(\tilde{\partial}c\phi^{*}_{1})]\Bigg{]}.$ (3)
@@
$S^{1}\ =\delta_{B}R^{1}\ =\ (R^{1},S^{0}),\qquad R^{1}=-\frac{1}{2}\int\ d^{D}x\ {\rm tr}\,\Big{[}\phi_{1}^{T}\tilde{G}(\phi\phi^{*}_{1})+(\tilde{\partial}c\phi_{1}^{*})^{T}\phi^{*}_{1}\Big{]}.$ (4)
However, whether the deformed total action $S(g)$ is free off-shell or not was left as an interesting future problem. In this letter, we tackle this issue and show that this higher-spin gauge model reduces to free off-shell as well.
@@
$\mathfrak{R}^{n}\ =\ \frac{1}{2}\int d^{D}x\ {\rm{tr}}\Bigg{[}\phi^{T}_{1}\Phi^{n}[\phi^{*}_{1}]+(\tilde{\partial}c\phi^{*}_{1})^{T}\Phi^{n-1}[\phi^{*}_{1}]\Bigg{]},$ (12)
where $\Phi^{0}[A]=A,\ \Phi[A]=\tilde{G}(\phi A),\ \Phi^{2}[A]=\tilde{G}(\phi\tilde{G}(\phi A))$, and so on. Noting that

### v2

@@ Abstract
By applying the BRST deformation scheme[5] in the BRST-antifield formalism, higher-spin gauge models of massless totally-symmetric tensors are constructed in [6]222 See [7, 8, 9, 10, 11] for other attempts by using BRST cohomological metod. . Subsequently, these are generalized to include massless totally-symmetric tensor-spinors in [12]. As an example, we consider the higher-spin gauge model of two bosons, a totally-symmetric rank-$n_{1}$ tensor $\phi^{\mu_{1}\cdots\mu_{n_{1}}}_{1}$ and a rank-$n$ tensor $\phi^{\mu_{1}\cdots\mu_{n}}$. We also introduce Grassmann-odd ghost fields, a rank-$(n_{1}-1)$ tensor $c_{1}^{\mu_{2}\cdots\mu_{n_{1}}}$ and a rank-$(n-1)$ tensor $c^{\mu_{2}\cdots\mu_{n}}$, and antifields $\{\phi^{*}_{1},\phi^{*},c_{1}^{*},c^{*}\}$ which have the same algebraic symmetries but opposite Grassmann parity. The action is given as
@@
where $g$ is the deformation parameter. $S^{0}$ represents the free action, while $S^{n}$ ($n\geq 1$) does the ($n+2$)-point vertex composed of $n+2$ (anti)fields, given by333 Hereafter, to simplify, we frequently omit totally-symmetric indices but this may not cause confusion.
$S^{0}\ =$ $\int d^{D}x\ \text{tr}\Bigg{[}\frac{1}{2}\phi_{1}^{T}G(\phi_{1})+\phi^{*}_{1}{}^{T}\partial c_{1}+\frac{1}{2}\phi^{T}G(\phi)+\phi^{*}{}^{T}\partial c\Bigg{]},$ (2)
$S^{n}\ =$ $\int d^{D}x\ \text{tr}\Bigg{[}\frac{1}{2}\phi^{T}_{1}\Phi^{n}[G(\phi_{1})]+(\tilde{\partial}c\phi^{*}_{1})^{T}\Phi^{n-1}[G(\phi_{1})]+\frac{1}{2}(\tilde{\partial}c\phi^{*}_{1})^{T}\Phi^{n-2}[\tilde{G}(\tilde{\partial}c\phi^{*}_{1})]\Bigg{]},$ (3)
where $\Phi^{0}[A]=A,\ \Phi[A]=\tilde{G}(\phi A),\ \Phi^{2}[A]=\tilde{G}(\phi\tilde{G}(\phi A))$, and so on. For an arbitrary totally-symmetric rank-$n$ tensor $A$, we have defined $\tilde{G}(A)$ as $\tilde{G}(A)\equiv G(A)+\frac{1}{2}\eta\partial\partial A^{\prime\prime}$ and $G(A)$ as $G(A)\equiv F(A)-\frac{1}{2}\eta F^{\prime}(A)$. $F$ is the Fronsdal tensor [13] defined by444 Our convention for symmetrization of indices is $\partial_{(\mu_{1}}\cdots\partial_{\mu_{r}}A_{\mu_{r+1}\cdots\mu_{n})\nu_{1}\cdots\nu_{m}}=\frac{1}{r!(n-r)!}\sum_{\{\mu_{1},\ldots,\mu_{n}\}}\partial_{\mu_{1}}\cdots\partial_{\mu_{r}}A_{\mu_{r+1}\cdots\mu_{n}\nu_{1}\cdots\nu_{m}},$ where $\{\mu_{1},\ldots,\mu_{n}\}$ indicates that the sum is taken over permutations of $\mu_{1},\ldots,\mu_{n}$.
$F_{\mu_{1}\cdots\mu_{n}}(A)=$ $\square A_{\mu_{1}\cdots\mu_{n}}-\partial_{(\mu_{1}}\partial\cdot A_{\mu_{2}\cdots\mu_{n})}+\partial_{(\mu_{1}}\partial_{\mu_{2}}A^{\prime}_{\mu_{3}\cdots\mu_{n})},$ (4)
where $\square=\eta^{\mu\nu}\partial_{\mu}\partial_{\nu}$ and $\eta^{\mu\nu}={\rm diag}(-1,+1,\ldots,+1)$. A prime on a tensor represents the trace, namely $A^{\prime}_{\mu_{3}\cdots\mu_{n}}=A_{\mu_{3}\cdots\mu_{n}}{}^{\rho}{}_{\rho}$ The divergence of $A$ is expressed as $\partial\cdot A_{\mu_{2}\cdots\mu_{n}}=\partial^{\mu_{1}}A_{\mu_{1}\mu_{2}\cdots\mu_{n}}$ We require that the ghost $c$ is traceless for $n\geq 3$, so that $F(\phi)$ is invariant under the BRST transformation $\delta_{B}\phi=\partial c$. For notational simplicity, we used a matrix notation in (2) and (3). For example, the integrand of the second term in (3) at $n=1$ means
~$\tilde{\partial}_{(\mu_{1}}c_{\mu_{2}\cdots\mu_{q})\rho_{1}\cdots\rho_{q}}{\phi^{*}}^{\rho_{1}\cdots\rho_{q}\nu_{1}\cdots\nu_{p}}_{1}G_{\nu_{1}\cdots\nu_{p}}{}^{\mu_{1}\cdots\mu_{q}}(\phi_{1}).$ (5)
Here $p+q=n_{1}$ and $2q=n$. The $\phi_{1}$ and ${\phi^{*}}_{1}$ denote $d(p)\times d(q)$ matrices.
@@
$S^{1}\ =\delta_{B}R^{1}\ =\ (R^{1},S^{0}),\qquad R^{1}=-\frac{1}{2}\int\ d^{D}x\ {\rm tr}\,\Big{[}\phi_{1}^{T}\tilde{G}(\phi\phi^{*}_{1})+(\tilde{\partial}c\phi_{1}^{*})^{T}\phi^{*}_{1}\Big{]},$ (6)
where the antibracket $(X,Y)$ for two functionals, $X[\Phi^{A},\Phi^{*}_{A}]$ and $Y[\Phi^{A},\Phi^{*}_{A}]$, is defined by
$(X,Y)\equiv X\frac{\overleftarrow{\delta}}{\delta\Phi^{A}}\frac{\overrightarrow{\delta}}{\delta\Phi^{*}_{A}}Y+X\frac{\overleftarrow{\delta}}{\delta\Phi^{*}_{A}}\frac{\overrightarrow{\delta}}{\delta\Phi^{A}}Y,$ (7)
where $\Phi^{*}_{A}$ are the antifields corresponding to $\Phi^{A}=\{\phi_{1},\phi,c_{1},c\}$. However, whether the deformed total action $S(g)$ is free off-shell or not was left as an interesting future problem. In this letter, we tackle this issue and show that this higher-spin gauge model reduces to free off-shell as well.
@@
$\mathfrak{R}^{n}\ =\ \frac{1}{2}\int d^{D}x\ {\rm{tr}}\Bigg{[}\phi^{T}_{1}\Phi^{n}[\phi^{*}_{1}]+(\tilde{\partial}c\phi^{*}_{1})^{T}\Phi^{n-1}[\phi^{*}_{1}]\Bigg{]}.$ (15)
Noting that